Orpheum Productions in Nevada is considering three mutually exclusive alternatives for lighting enhancements to one of its recording studios. Each enhancement will increase revenues by attracting directors who prefer this lighting style. The cash flow details, in thousands of dollars, for these enhancements are shown in the chart below. MARR is 10%/year.

End of Year

Light Bar

Sliding Spots

Reflected Beam

0 -$10,000 -$18,000 -$27,000
1 $2,000 $3,500 $0
2 $2,000 $3,500 $2,700
3 $2,000 $3,500 $5,400
4 $2,000 $3,500 $8,100
5 $2,000 $3,500 $10,800
6 $2,000 $3,500

$13,500

Show the annual worth for each option
Answer #1

MARR = 10%

AW of light bar = -10000*(A/P,10%,6) + 2000 = -10000*0.229607 + 2000 = -296.07 thousands

AW of Sliding spots = -18000*(A/P,10%,6) + 3500 = -18000*0.229607 + 3500 = -623.926 thousands

AW of reflected beam = -27000*(A/P,10%,6) + 2700*(A/G,10%,6)

= -27000*0.229607 + 2700* 2.22355 = -195.804 thousands